After this replacement, we measured the spectra of some signals related to PMC (error signal, feedback signal, trans/reflect power, etc,. ) in order to check if there is any change before and after the power supply is replaced. Fig. 1 shows the results. The black line is the spectrum before the replacement (2021/01/07 00:35 UTC), the blue one is the spectrum after the replacement (02:38 UTC). As you can see, the blue one has a spike at 6 kHz.
We found that the overall gain (K1:PSL-PMC_COMMON_GAIN) gets larger when PMC locks again. (14 dB -> 22 dB, fig. 2). So we considered that the control is about to oscillate due to the gain increase, I reduced the gain 22dB to 14 dB. then the spikes are disappeared in red lines.
What we are wondering is why the gain was so large this time..